OpenStudy (anonymous):

the area of a rectange is 35 square feet. the length is two feet longer than the width. what is the width of the rectangle?

OpenStudy (anonymous):

its quadratic formula

OpenStudy (callisto):

let w be the width, length = w+2 Area of rectangle = length x width 35 = w ( w+2) 35 = w^2 + 2w w^2 + 2w - 35 =0 w^2 -5x +7w -35 =0 w(w-5) + 7(w-5) =0 (w+7)(w-5) =0 w+7 =0 or w-5 =0 Can you solve the equations now? Reject the answer of negative value as length cannot be of negative value...